OpenStudy (anonymous):

help! Solve for x: -3(x + 3) = -3(x + 1) - 5.

OpenStudy (therealmeeeee):

i will

OpenStudy (anonymous):

these are the options 6 -6 All real numbers No solution

OpenStudy (therealmeeeee):

x^5 - x - 7 = 0 f(x) = x^5 - x - 7 f'(x) = 5x^4 - 1 x2 = x1 - (x1^5 - x1 - 7)/(5x1^4 - 1) = 1 + 7/4 = 11/4 = 2.75 x3 = 162843/72949 = 0.86146

OpenStudy (anonymous):

can u help @TheRealMeeeee

OpenStudy (therealmeeeee):

wit wat

OpenStudy (therealmeeeee):

with wat @lulu22

OpenStudy (anonymous):

the question at th top @TheRealMeeeee

OpenStudy (therealmeeeee):

i answered it already

OpenStudy (anonymous):

i dont see it @the

OpenStudy (therealmeeeee):

x^5 - x - 7 = 0 f(x) = x^5 - x - 7 f'(x) = 5x^4 - 1 x2 = x1 - (x1^5 - x1 - 7)/(5x1^4 - 1) = 1 + 7/4 = 11/4 = 2.75 x3 = 162843/72949 = 0.86146

OpenStudy (anonymous):

so the answer is what @TheRealMeeeee

OpenStudy (therealmeeeee):

0.86146

OpenStudy (anonymous):

that is not a choice @TheRealMeeeee

OpenStudy (therealmeeeee):

wat thats wat i got then idk

OpenStudy (therealmeeeee):

thats wat i got

OpenStudy (anonymous):

okk ught thanks anyways @TheRealMeeeee

OpenStudy (therealmeeeee):

np beautiful

OpenStudy (anonymous):

6 -6 All real numbers No solution

OpenStudy (anonymous):

those are the choices tho

OpenStudy (anonymous):

-3(x + 3) = -3(x + 1) - 5 -3x - 9 = -3x - 3 - 5 -3x + 3x = -3 -5 + 9 0 = 1 (incorrect) NO SOLUTION

OpenStudy (anonymous):

okkk thank u soo much @kelliegirl33

OpenStudy (anonymous):

anytime :)